# N Hale, E Thomann, JAC Weideman

### v1

@@ 4 Numerical method
Motivated by the semi-infinite domain, the smoothness and exponential decay of the solutions, and a desire for high accuracy, we solve (3.1) numerically using a Laguerre spectral collocation method with scaling appropriate to match the $v(t)={\rm O}(e^{-t})$ behaviour for large $t$. To improve efficiency we follow (Mastroianni and Monegato, 2008) in using a truncated Laguerre method. In particular, we approximate the solution on the first $\lceil M\sqrt{N}\rceil$ points of an $N$-point Laguerre grid, with $M$ determined heuristically based on experiments below. The functional terms, $v(\alpha t)$ and $v^{2}(\alpha t)$, in the equation are implemented with the barycentric resampling approach described in (Hale, 2023a). In short, this approach uses the barycentric interpolation formula to construct a matrix, $P_{\alpha}$, that interpolates the weighted polynomial representation of the solution, $v$, evaluated at the vector of truncated Laguerre points, $\bm{t}$, to the dilated points, $\alpha\bm{t}$, so that $v(\alpha\bm{t})\approx P_{\alpha}v(\bm{t})$. This yields the nonlinear system of equations,
$(D+I-2P_{\alpha})\bm{v}-P_{\alpha}(\bm{v}\circ\bm{v})=0,$ (4.1)
@@
We begin by verifying the efficacy of the Laguerre discretisation by solving the linear problem (2.1) and comparing to the series solution (2.2). To do so, we omit the nonlinear term from (4.1) and obtain nontrivial solutions by computing the nullspace of the matrix $D+I-2P_{\alpha}$. Figure 4 shows the error between the spectral collocation and series solutions for various $\alpha_{n}$ as the number of degrees of freedom in the collocation method is increased. The error curves indicate rapid convergence up to a point where the effects of a sub-optimal truncation point degrades the quality of the solution. Rather than trying to find the optimal scaling and discretisation size for each $\alpha$, based on the results in Figure 4, we take $M=6$ and $N=700$ (so that $\lceil M\sqrt{N}\rceil=159$) in our remaining experiments to provide a balance between accuracy and execution time.
Figure 4: Errors in the truncated Laguerre spectral discretisation of (2.1) for the characteristic solutions at $\alpha=2,2^{1/3},$ and $2^{1/5}$, normalized such that $\|u\|_{2}=1$. Shown are the maximum grid errors as compared to the series solution (2.2), computed with sufficiently many terms. The series solution is computed in extended precision to mitigate cancellation errors introduced by the large transient growth in its terms for large $n$.
Having verified that the Laguerre discretisation provides accurate solutions to the linear problem (2.1), we now use it to solve the nonlinear problem (3.1). We first do so in the vicinity of the characteristic values, $\alpha_{n}=2^{1/n}$. In particular, we use the perturbation approximations derived in section 3 as initial guesses which are then refined using a Newton iteration applied to (4.1). Some examples have already been shown in Figure 3. For each $n$, these solutions are then used to seed a pseudo-arclength continuation, as described by (Farrell et al., 2015) and implemented in Chebfun (Driscoll et al., 2014), to compute solutions over a desired range of $\alpha$.
A MATLAB code that implements the method described here is available at (Hale, 2023b). It can be used to reproduce all figures of this paper.
@@ 5 A family of solutions
Using the numerical procedure described in the previous section, we now explore the family of solutions in class (A) as defined in the first paragraph of section 3, i.e., the solutions to  (1.2) that satisfy $u(0)=1$ and $u(t)=1+{\rm O}(e^{-t})$ as $t\to\infty$. Our aim is to catalog for each value of $\alpha$ the number of possible solutions in this class. Recall that for each value of $\alpha$ the constant solution $u=1$ is in this class, so we focus on counting the additional solutions. Excluded in the counting is the solution that satisfies $u(t)={\rm O}(e^{-t})$ as $t\to\infty$, as it is not in class (A). Possible solutions that obey an asymptotic behaviour different from ${\rm O}(e^{-t})$ are excluded in the counting as well (but, as mentioned previously, such solutions are unlikely to be picked up by the numerical method used here).
A bifurcation diagram for solutions in this family is shown in the first frame of Figure 5, and solutions to (1.2) are shown in the other frames. The bifurcation diagram shows the values of the $L^{2}$-norm (squared) of solutions $v$ to (3.1) as $\alpha$ is varied. Points at which the curves touch the horizontal axis correspond to $v=0$, therefore representing the constant solution $u=1$. This happens at the critical values of $\alpha$, namely $\alpha_{n}=2^{1/n}$, $n=1,2,\ldots$. In the vicinity of these points, the continuation procedure described in the previous section was used to generate further solutions.
Figure 5: The first frame shows the bifurcation diagram and the remaining frames show solutions to (1.2) various values of $\alpha$. The black dotted solution decaying to zero is the one in class (B) as defined in the first paragraph of section 3. The multicoloured solutions are in class (A). Note the labeling: the letters at the top of the bifurcation diagram correspond to the various solution plots. Dotted (resp. solid) curves are used to represent solutions whose first turning point is a local maximum (resp. minimum). (An animated version of the figure can be seen at https://youtu.be/aBTvLujm65c.)
@@
For more detail about these solutions, consider Figure 6, which zooms in on the interval around $n=2$ in the bifurcation diagram. When $n$ is less than $2$ by a sufficient margin, the only solution in class (A) (other than the constant solution) is a minus-solution such as the one shown with the label H. (This solution is similar to the one computed in (Dascaliuc et al., 2023, Fig. 2) for a slightly different value of $\alpha$.) The solution marked I corresponds to the critical value of $n\approx 1.91$ ($\alpha\approx 1.44$) where a first additional solution is generated. With increasing $n$ this solution bifurcates into two plus-solutions as shown in the subplot labeled J. When $n$ reaches the value $2$, one of these plus-solutions (the one with the smaller norm) momentarily collapses to the constant solution at the bifurcation point (subplot K) and then re-emerges as a minus-solution for larger $n$ (subplot L). Observe that the two solutions of small norm shown in subplots J and L in Figure 6 correspond to the two solutions shown in the middle frame of Figure 3. These are therefore examples of the additional solutions alluded to at the end of section 3.
@@
Returning to Figure 5 and continuing to increase the value of $n$, one notices a similar pattern. In each interval $(n-\delta,n)$, with $2\leq n\leq 6$ and $0<\delta\ll 1$, two new solutions are generated. When created, both of these are plus-solutions but one of them switches to a minus-solution once the value of $n$ is exceeded. The magnitude of $\delta$ is about $0.09$ when $n=2$ but then decreases rapidly with $n$. (We conjecture that the inequality $2\leq n\leq 6$ can be replaced by $2\leq n<\infty$. But numerical experimentation in the range $n\geq 6.5$ is complicated by the fact that the perturbation approximation (3.7) becomes less effective as a starting guess, and therefore non-convergence of the Newton iteration becomes an issue.)
@@ 6 Conclusions
The methodology presented here for solving (1.2) has several ingredients: a Laguerre spectral collocation method for solving the equation numerically, a perturbation method for generating initial guesses for a Newton iteration, and a continuation method for generating further solutions. These techniques are well-established for standard ordinary differential equations, but we believe their combination to solve a nonlinear functional differential equation is new.
@@
The methodology should also enable one to consider the following generalization of (1.2),
$u^{\prime}(t)+au(t)=bu^{m}(\alpha t),$ (6.1)
@@
Another avenue of investigation is how to solve (1.2) for other initial conditions. One possibility might be to use the $u(0)=1$ solutions in combination with continuation on the $u_{0}$ parameter to generate solutions corresponding to nearby initial conditions.
@@ Appendix A Appendix
By multiplying (3.1) by $t^{j}$ and integrating by parts, it follows that [^3] Here, and below, moments with negative indices are included for ease of notation. They should be taken to be identically zero.
$(\alpha^{j+1}-2)\mu_{j}(v)-j\alpha^{j+1}\mu_{j-1}(v)-\mu_{j}(v^{2})=0,$ (A.2)
@@
Now, for $n=1,2,\ldots$, consider the perturbation $\alpha=\alpha_{n}+\epsilon$, where $\alpha_{n}=2^{1/n}$, and the expansion $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$, as in section 3. As before, $v_{0}$ is the solution to (2.1) with $\alpha=\alpha_{n}$. Substitute this into (A.2) and use the binomial theorem to expand $(\alpha_{n}+\epsilon)^{j+1}$. Equating the first two powers of $\epsilon$ gives
$(\alpha^{j+1}_{n}-2)\mu_{j}(v_{0})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{0})=0,$ (A.3)
@@
$(\alpha^{j+1}_{n}-2)\mu_{j}(v_{1})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{1})+(j+1)\alpha_{n}^{j}\mu_{j}(v_{0})-j(j+1)\alpha_{n}^{j}\mu_{j-1}(v_{0})-\mu_{j}(v_{0}^{2})=0,$ (A.4)
@@
$(\alpha^{j+1}_{n}-2)\mu_{j}(v_{1})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{1})+\frac{2(j+1)}{\alpha}\mu_{j}(v_{0})-\mu_{j}(v_{0}^{2})=0.$ (A.5)
At this stage it is useful to note that, because $v_{0}$ is known—at least, up to the scaling constants $C_{n}$ that we are presently trying to determine—all moments involving $v_{0}$ can be computed numerically; see (A.19) below. This is not the case for moments of $v_{1}$, however, since $v_{1}$ is unknown, and hence the purpose in the following is to eliminate such moments.
With $j=n-1$, we have $\alpha_{n}^{j+1}=2$ and equations (A.3) and (A.5) simplify to
$(n-1)\mu_{n-2}(v_{0})=0,$ (A.6)
@@
$-2(n-1)\mu_{n-2}(v_{1})+\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})-\mu_{n-1}(v_{0}^{2})=0,$ (A.7)
respectively. The fact that the $\mu_{n-1}(v_{1})$ term is eliminated in the latter expression is precisely the reason for considering the $(n-1)$th moment. This will allow us to develop a telescoping recurrence in order to eliminate moments involving the unknown $v_{1}$. To this end, note that (A.7) can be rearranged as
$\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\mu_{n-1}(v_{0}^{2})+2(n-1)\mu_{n-2}(v_{1}),$ (A.8)
@@
Now consider $j<n-1$. It follows from (A.3) and (A.6) that
$\mu_{j}(v_{0})=0,\qquad j<n-1.$ (A.9)
This causes the third term on the left of (A.5) to vanish, which gives the recurrence
$\mu_{j}(v_{1})=\frac{1}{\alpha_{{n}}^{j+1}-2}\left(\mu_{j}(v_{0}^{2})+j\alpha_{{n}}^{j+1}\mu_{j-1}(v_{1})\right),\qquad j<n-1.$ (A.10)
Returning now to (A.8) and applying (A.10) with $j=n-2$ gives
$\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\mu_{n-1}(v_{0}^{2})+\frac{2(n-1)}{\alpha^{n-1}_{n}-2}\left(\mu_{n-2}(v_{0}^{2})+(n-2)\alpha_{n}^{n-1}\mu_{n-3}(v_{1})\right).$ (A.11)
@@
$2\sqrt{2}\,\mu_{1}(v_{0})=\mu_{1}(v_{0}^{2})-(2+\sqrt{2})\,\mu_{0}(v_{0}^{2}).$ (A.12)
If $n>2$ then further substitutions of (A.10) into (A.11) are made until the resulting series likewise terminates with the vanishing of the quantity involving $v_{1}$ and only moments of $v_{0}$ and $v_{0}^{2}$ remain. The general formula involves the following linear combination of the $\mu_{j}(v_{0}^{2})$ on the right
$\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\sum_{j=0}^{n-1}w_{j}\mu_{j}(v_{0}^{2}),\qquad n=1,2,\ldots,$ (A.13)
@@
$w_{j-1}=\frac{j\alpha^{j+1}_{n}}{\alpha^{j}_{n}-2}w_{j},\quad j=n-1,\ldots,1.$ (A.14)
Letting $v_{0}(t)=C_{n}E(t;\alpha_{n})$ as in (2.2), it follows that (A.13) becomes
$\frac{2n}{\alpha_{n}}C_{n}\mu_{n-1}(E_{n})=C_{n}^{2}\sum_{j=0}^{n-1}w_{j}\mu_{j}(E^{2}_{n}),$ (A.15)
@@
$C_{n}=\frac{2n\mu_{n-1}(E_{n})}{\alpha_{n}\sum_{j=0}^{n-1}w_{j}\mu_{j}(E^{2}_{n})},\quad n=1,2,\ldots.$ (A.16)
@@
$C_{1}=\frac{\mu_{0}(E_{1})}{\mu_{0}(E_{1}^{2})},$ (A.17)
a formula already given in (3.6). For $n=2$ ($\alpha_{2}=\sqrt{2}$) the formula is
$C_{2}=\frac{2\sqrt{2}\,\mu_{1}(E_{2})}{\mu_{1}(E^{2}_{2})-(2+\sqrt{2})\mu_{0}(E_{2}^{2})},$ (A.18)
which also follows directly from (A.12). For larger values of $n$ the formulas are not as clean as the surds do not cancel elegantly.
Explicit series formulas for the moments in (A.16)–(A.18) can be obtained by integrating the series (2.2) term-by-term. In particular,
$\mu_{j}(E_{n})=j!\sum_{k=0}^{\infty}\frac{b_{k}}{\alpha_{n}^{(j+1)k}}\qquad\text{and}\qquad\mu_{j}(E^{2}_{n})=j!\sum_{k=0}^{\infty}\sum_{\ell=0}^{\infty}\frac{b_{k}b_{\ell}}{(\alpha_{n}^{k}+\alpha_{n}^{\ell})^{j+1}},$ (A.19)
@@
$b_{k}=\left\{\begin{array}[]{cl}1,&k=0,\\ \displaystyle\frac{2^{k}}{(1-\alpha_{n})\cdots(1-\alpha_{n}^{k})},&k>0.\end{array}\right.$ (A.20)
Although the coefficients in these series have exponential asymptotic decay, they have alternating signs and large transient growth, making them numerically unstable, particularly when $n$ is large. Extended precision is therefore required to compute the moments accurately from these formulas. Alternatively, one can approximate the moments with numerical integration, but care is still required in evaluating (2.2).

### v2

@@ 4 Numerical method
Motivated by the semi-infinite domain, the smoothness and exponential decay of the solutions, and a desire for high accuracy, we solve (7) numerically using a Laguerre spectral collocation method with scaling appropriate to match the $v(t)={\rm O}(e^{-t})$ behaviour for large $t$. To improve efficiency we follow [17] in using a truncated Laguerre method. In particular, we approximate the solution on the first $\lceil M\sqrt{N}\rceil$ points of an $N$-point Laguerre grid, with $M$ determined heuristically based on experiments below. The functional terms, $v(\alpha t)$ and $v^{2}(\alpha t)$, in the equation are implemented with the barycentric resampling approach described in [14]. In short, this approach uses the barycentric interpolation formula to construct a matrix, $P_{\alpha}$, that interpolates the weighted polynomial representation of the solution, $v$, evaluated at the vector of truncated Laguerre points, $\bm{t}$, to the dilated points, $\alpha\bm{t}$, so that $v(\alpha\bm{t})\approx P_{\alpha}v(\bm{t})$. This yields the nonlinear system of equations,
(16) $(D+I-2P_{\alpha})\bm{v}-P_{\alpha}(\bm{v}\circ\bm{v})=0,$
@@
We begin by verifying the efficacy of the Laguerre discretisation by solving the linear problem (3) and comparing to the series solution (4). To do so, we omit the nonlinear term from (16) and obtain nontrivial solutions by computing the nullspace of the matrix $D+I-2P_{\alpha}$. Figure 4 shows the error between the spectral collocation and series solutions for various $\alpha_{n}$ as the number of degrees of freedom in the collocation method is increased. The error curves indicate rapid convergence up to a point where the effects of a sub-optimal truncation point degrades the quality of the solution. Rather than trying to find the optimal scaling and discretisation size for each $\alpha$, based on the results in Figure 4, we take $M=6$ and $N=700$ (so that $\lceil M\sqrt{N}\rceil=159$) in our remaining experiments to provide a balance between accuracy and execution time.
Figure 4: Errors in the truncated Laguerre spectral discretisation of (3) for the characteristic solutions at $\alpha=2,2^{1/3},$ and $2^{1/5}$, normalized such that $\|u\|_{2}=1$. Shown are the maximum grid errors as compared to the series solution (4), computed with sufficiently many terms. The series solution is computed in extended precision to mitigate cancellation errors introduced by the large transient growth in its terms for large $n$.
Having verified that the Laguerre discretisation provides accurate solutions to the linear problem (3), we now use it to solve the nonlinear problem (7). We first do so in the vicinity of the characteristic values, $\alpha_{n}=2^{1/n}$. In particular, we use the perturbation approximations derived in section 3 as initial guesses which are then refined using a Newton iteration applied to (16). Some examples have already been shown in Figure 3. For each $n$, these solutions are then used to seed a pseudo-arclength continuation, as described by [9] and implemented in Chebfun [8], to compute solutions over a desired range of $\alpha$.
A MATLAB code that implements the method described here is available at [13]. It can be used to reproduce all figures of this paper.
@@ 5 A family of solutions
Using the numerical procedure described in the previous section, we now explore the family of solutions in class (A) as defined in the first paragraph of section 3, i.e., the solutions to (2) that satisfy $u(0)=1$ and $u(t)=1+{\rm O}(e^{-t})$ as $t\to\infty$. Our aim is to catalog for each value of $\alpha$ the number of possible solutions in this class. Recall that for each value of $\alpha$ the constant solution $u=1$ is in this class, so we focus on counting the additional solutions. Excluded in the counting is the solution that satisfies $u(t)={\rm O}(e^{-t})$ as $t\to\infty$, as it is not in class (A). Possible solutions that obey an asymptotic behaviour different from ${\rm O}(e^{-t})$ are excluded in the counting as well (but, as mentioned previously, such solutions are unlikely to be picked up by the numerical method used here).
A bifurcation diagram for solutions in this family is shown in the first frame of Figure 5, and solutions to (2) are shown in the other frames. The bifurcation diagram shows the values of the $L^{2}$-norm (squared) of solutions $v$ to (7) as $\alpha$ is varied. Points at which the curves touch the horizontal axis correspond to $v=0$, therefore representing the constant solution $u=1$. This happens at the critical values of $\alpha$, namely $\alpha_{n}=2^{1/n}$, $n=1,2,\ldots$. In the vicinity of these points, the continuation procedure described in the previous section was used to generate further solutions.
Figure 5: The first frame shows the bifurcation diagram and the remaining frames show solutions to (2) various values of $\alpha$. The black dotted solution decaying to zero is the one in class (B) as defined in the first paragraph of section 3. The multicoloured solutions are in class (A). Note the labeling: the letters at the top of the bifurcation diagram correspond to the various solution plots. Dotted (resp. solid) curves are used to represent solutions whose first turning point is a local maximum (resp. minimum). (An animated version of the figure can be seen at https://youtu.be/aBTvLujm65c.)
@@
For more detail about these solutions, consider Figure 6, which zooms in on the interval around $n=2$ in the bifurcation diagram. When $n$ is less than $2$ by a sufficient margin, the only solution in class (A) (other than the constant solution) is a minus-solution such as the one shown with the label H. (This solution is similar to the one computed in [5, Fig. 2] for a slightly different value of $\alpha$.) The solution marked I corresponds to the critical value of $n\approx 1.91$ ($\alpha\approx 1.44$) where a first additional solution is generated. With increasing $n$ this solution bifurcates into two plus-solutions as shown in the subplot labeled J. When $n$ reaches the value $2$, one of these plus-solutions (the one with the smaller norm) momentarily collapses to the constant solution at the bifurcation point (subplot K) and then re-emerges as a minus-solution for larger $n$ (subplot L). Observe that the two solutions of small norm shown in subplots J and L in Figure 6 correspond to the two solutions shown in the middle frame of Figure 3. These are therefore examples of the additional solutions alluded to at the end of section 3.
@@
Returning to Figure 5 and continuing to increase the value of $n$, one notices a similar pattern. In each interval $(n-\delta,n)$, with $2\leq n\leq 6$ and $0<\delta\ll 1$, two new solutions are generated. When created, both of these are plus-solutions but one of them switches to a minus-solution once the value of $n$ is exceeded. The magnitude of $\delta$ is about $0.09$ when $n=2$ but then decreases rapidly with $n$. (We conjecture that the inequality $2\leq n\leq 6$ can be replaced by $2\leq n<\infty$. But numerical experimentation in the range $n\geq 6.5$ is complicated by the fact that the perturbation approximation (13) becomes less effective as a starting guess, and therefore non-convergence of the Newton iteration becomes an issue.)
@@ 6 Conclusions
The methodology presented here for solving (2) has several ingredients: a Laguerre spectral collocation method for solving the equation numerically, a perturbation method for generating initial guesses for a Newton iteration, and a continuation method for generating further solutions. These techniques are well-established for standard ordinary differential equations, but we believe their combination to solve a nonlinear functional differential equation is new.
@@
The methodology should also enable one to consider the following generalization of (2),
(17) $u^{\prime}(t)+au(t)=bu^{m}(\alpha t),$
@@
Another avenue of investigation is how to solve (2) for other initial conditions. One possibility might be to use the $u(0)=1$ solutions in combination with continuation on the $u_{0}$ parameter to generate solutions corresponding to nearby initial conditions.
@@ Appendix A Appendix
By multiplying (7) by $t^{j}$ and integrating by parts, it follows that [^3] Here, and below, moments with negative indices are included for ease of notation. They should be taken to be identically zero.
(19) $(\alpha^{j+1}-2)\mu_{j}(v)-j\alpha^{j+1}\mu_{j-1}(v)-\mu_{j}(v^{2})=0,$
@@
Now, for $n=1,2,\ldots$, consider the perturbation $\alpha=\alpha_{n}+\epsilon$, where $\alpha_{n}=2^{1/n}$, and the expansion $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$, as in section 3. As before, $v_{0}$ is the solution to (3) with $\alpha=\alpha_{n}$. Substitute this into (19) and use the binomial theorem to expand $(\alpha_{n}+\epsilon)^{j+1}$. Equating the first two powers of $\epsilon$ gives
(20) $(\alpha^{j+1}_{n}-2)\mu_{j}(v_{0})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{0})=0,$
@@
(21) $(\alpha^{j+1}_{n}-2)\mu_{j}(v_{1})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{1})+(j+1)\alpha_{n}^{j}\mu_{j}(v_{0})-j(j+1)\alpha_{n}^{j}\mu_{j-1}(v_{0})-\mu_{j}(v_{0}^{2})=0,$
@@
(22) $(\alpha^{j+1}_{n}-2)\mu_{j}(v_{1})-j\alpha^{j+1}_{n}\mu_{j-1}(v_{1})+\frac{2(j+1)}{\alpha}\mu_{j}(v_{0})-\mu_{j}(v_{0}^{2})=0.$
At this stage it is useful to note that, because $v_{0}$ is known—at least, up to the scaling constants $C_{n}$ that we are presently trying to determine—all moments involving $v_{0}$ can be computed numerically; see (36) below. This is not the case for moments of $v_{1}$, however, since $v_{1}$ is unknown, and hence the purpose in the following is to eliminate such moments.
With $j=n-1$, we have $\alpha_{n}^{j+1}=2$ and equations (20) and (22) simplify to
[…]
@@
(24) $-2(n-1)\mu_{n-2}(v_{1})+\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})-\mu_{n-1}(v_{0}^{2})=0,$
respectively. The fact that the $\mu_{n-1}(v_{1})$ term is eliminated in the latter expression is precisely the reason for considering the $(n-1)$th moment. This will allow us to develop a telescoping recurrence in order to eliminate moments involving the unknown $v_{1}$. To this end, note that (24) can be rearranged as
(25) $\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\mu_{n-1}(v_{0}^{2})+2(n-1)\mu_{n-2}(v_{1}),$
@@
Now consider $j<n-1$. It follows from (20) and (23) that
(26) $\mu_{j}(v_{0})=0,\qquad j<n-1.$
This causes the third term on the left of (22) to vanish, which gives the recurrence
(27) $\mu_{j}(v_{1})=\frac{1}{\alpha_{{n}}^{j+1}-2}\left(\mu_{j}(v_{0}^{2})+j\alpha_{{n}}^{j+1}\mu_{j-1}(v_{1})\right),\qquad j<n-1.$
Returning now to (25) and applying (27) with $j=n-2$ gives
(28) $\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\mu_{n-1}(v_{0}^{2})+\frac{2(n-1)}{\alpha^{n-1}_{n}-2}\left(\mu_{n-2}(v_{0}^{2})+(n-2)\alpha_{n}^{n-1}\mu_{n-3}(v_{1})\right).$
@@
(29) $2\sqrt{2}\,\mu_{1}(v_{0})=\mu_{1}(v_{0}^{2})-(2+\sqrt{2})\,\mu_{0}(v_{0}^{2}).$
If $n>2$ then further substitutions of (27) into (28) are made until the resulting series likewise terminates with the vanishing of the quantity involving $v_{1}$ and only moments of $v_{0}$ and $v_{0}^{2}$ remain. The general formula involves the following linear combination of the $\mu_{j}(v_{0}^{2})$ on the right
(30) $\frac{2n}{\alpha_{n}}\mu_{n-1}(v_{0})=\sum_{j=0}^{n-1}w_{j}\mu_{j}(v_{0}^{2}),\qquad n=1,2,\ldots,$
@@
(31) $w_{j-1}=\frac{j\alpha^{j+1}_{n}}{\alpha^{j}_{n}-2}w_{j},\quad j=n-1,\ldots,1.$
Letting $v_{0}(t)=C_{n}E(t;\alpha_{n})$ as in (4), it follows that (30) becomes
(32) $\frac{2n}{\alpha_{n}}C_{n}\mu_{n-1}(E_{n})=C_{n}^{2}\sum_{j=0}^{n-1}w_{j}\mu_{j}(E^{2}_{n}),$
@@
(33) $C_{n}=\frac{2n\mu_{n-1}(E_{n})}{\alpha_{n}\displaystyle\sum_{j=0}^{n-1}w_{j}\mu_{j}(E^{2}_{n})},\quad n=1,2,\ldots.$
@@
(34) $C_{1}=\frac{\mu_{0}(E_{1})}{\mu_{0}(E_{1}^{2})},$
a formula already given in (12). For $n=2$ ($\alpha_{2}=\sqrt{2}$) the formula is
(35) $C_{2}=\frac{2\sqrt{2}\,\mu_{1}(E_{2})}{\mu_{1}(E^{2}_{2})-(2+\sqrt{2})\mu_{0}(E_{2}^{2})},$
which also follows directly from (29). For larger values of $n$ the formulas are not as clean as the surds do not cancel elegantly.
Explicit series formulas for the moments in (33)–(35) can be obtained by integrating the series (4) term-by-term. In particular,
(36) $\mu_{j}(E_{n})=j!\sum_{k=0}^{\infty}\frac{b_{k}}{\alpha_{n}^{(j+1)k}}\qquad\text{and}\qquad\mu_{j}(E^{2}_{n})=j!\sum_{k=0}^{\infty}\sum_{\ell=0}^{\infty}\frac{b_{k}b_{\ell}}{(\alpha_{n}^{k}+\alpha_{n}^{\ell})^{j+1}},$
@@
(37) $b_{k}=\left\{\begin{array}[]{cl}1,&k=0,\\ \displaystyle\frac{2^{k}}{(1-\alpha_{n})\cdots(1-\alpha_{n}^{k})},&k>0.\end{array}\right.$
Although the coefficients in these series have exponential asymptotic decay, they have alternating signs and large transient growth, making them numerically unstable, particularly when $n$ is large. Extended precision is therefore required to compute the moments accurately from these formulas. Alternatively, one can approximate the moments with numerical integration, but care is still required in evaluating (4).